OpenStudy (anonymous):

what number results when the square of 14 is reduced by 56 and then divided by the additive inverse of (-2*7)

OpenStudy (anonymous):

\[\frac{14^2 -56}{-1\cdot(- 2 \cdot 7)} = \frac{196 - 56}{14}=\frac{140}{14}=10\]